Find three consecutive integers such that the square of the first increased by twice the second is 3 less than four times the third.
This is a quadratic equation

Answer by mananth(16949)

let the integers be x,x+1,x+2
x^2+2(x+1)=4(x+2)-3
x^2+2x+2=4x+8-3
x^2-2x-3=0
x^2-3x+x-3=0
x(x-3)+1(x-3)=0
(x-3)(x+1)=0
x= 3 OR -1
The integers are 3,4,5
OR
-1,-2,-3